OpenStudy (anonymous):

Calculate the double integral ∫∫R (10x + 2y + 20 ) dA where R is the region: 0 ≤ x ≤ 1, 0 ≤ y ≤ 5.

OpenStudy (turingtest):

\[\int\limits \int\limits10x+2y+20dA=\int\limits_{0}^{1}\int\limits_{0}^{5}10x-2y+20dydx\]\[=\int\limits_{0}^{1}10x(5)(5)^2+20(5)dx=\int\limits_{0}^{1}50xdx=25(1)^2=25\]

OpenStudy (turingtest):

sorry los a + sign on the second line, should be 10x(5)+(5)^2+20(5)

OpenStudy (anonymous):

yeah i was gonna say that !

OpenStudy (turingtest):

good, then you got it?

OpenStudy (anonymous):

yeah so its 175 !?

OpenStudy (turingtest):

wait a minute let me take it from where I messed up...

OpenStudy (anonymous):

ok

OpenStudy (turingtest):

argh! it's 150 stupid thing, want me to show it from the top?

OpenStudy (turingtest):

the program can be stupid I mean...

OpenStudy (anonymous):

ok cause now i'm lost !

OpenStudy (anonymous):

i understand !

OpenStudy (turingtest):

\[\int\limits_{0}^{1}\int\limits_{0}^{5}10x+2y+20dydx=\int\limits_{0}^{1}10x(5)+5^2+20(5)dx\]\[\int\limits_{0}^{1}50x+125dx=25+125=50\]

OpenStudy (turingtest):

see that, the last line is clearly 25+125=150 there is a delay, sorry for the confusion.

OpenStudy (anonymous):

yeah thanks ! i actually did it myself but thanks for all the explanations !